What is the probability that a roll of 2 number cubes roll a sum of 24?
If they are standard cubes - with numbers from 1 to 6 - the probability is 0.
What is the difference between a fraternal twins and identical twins?
If two (or more) eggs are fertilised and develop together you will get fraternal twins. Identical twins are the result of a single fertilised egg dividing into two embryos.
What happens after arthroplasty?
Patients who have undergone arthroplasty must be careful not to over stress or destabilize the joint
How do you find p(b) when p(a) is 23 p(ba) is 12 and p(a U b) is 45 and is a dependent event?
There are symbols missing from your question which I cam struggling to guess and re-insert.
p(a) = 2/3
p(b ??? a) = 1/2
p(a ∪ b) = 4/5
p(b) = ?
Why use the set notation of Union on the third given probability whereas the second probability has something missing but the "sets" are in the other order, and the order wouldn't matter in sets.
There are two possibilities:
1) The second probability is: p(b ∩ a) = p(a ∩ b) = 1/2
→ p(a) + p(b) = p(a ∪ b) + p(a ∩ b)
→ p(b) = p(a ∪ b) + p(a ∩ b) - p(a)
= 4/5 + 1/2 - 2/3
= 24/30 + 15/30 - 20/30
= 19/30
2) The second and third probabilities are probabilities of "given that", ie:
p(b|a) = 1/2
p(a|b) = 4/5
→ Use Bayes theorem:
p(b)p(a|b) = p(a)p(b|a)
→ p(b) = (p(a)p(b|a))/p(a|b)
= (2/3 × 1/2) / (4/5)
= 2/3 × 1/2 × 5/4 = 5/12
What are the strategies for finding theoretical probabilities?
I have no clue I'm wondering the same thing!
Joint probability is the probability that two or more specific outcomes will occur in an event. An example of joint probability would be rolling a 2 and a 5 using two different dice.
If Two dice are rolled Find the probability of getting adouble?
One cube can land in 6 different ways.
Two cubes can land in (6 x 6) = 36 different ways.
Doubles can come up in 6 different ways.
Probability of a double = (6/36) = 1/6 = 16-2/3 %
What is the probability of getting heads on all 2 tosses if a fair coin is tossed 5 times?
With 5 coin tosses there are 32 possible outcomes. 10 of these have exactly 2 heads, and 26 of these have 2 or more heads.
What is the probability with 2 dices getting roles on the same number?
The probability of rolling two 6-sided dice and getting the same number on each is 1/32, or "one out of thirty two".
This is because there are 6 oppurtunities on each dice, and the probability of getting the same of the 6 oppurtunites on two dice is 6 squared, which is 32.
not likely and near impossible
What is the relationship between probability and the binomial theorem?
What is the symbol for a Probability of success in a binomial trial?
How is math used for a Chess Player?
I think the only way math is used in chess is to determine the point values of the pieces.
This thinking is common among newer players. But point value certainly isn't everything. True, math itself is not the biggest thing needed to be good at chess, but it does play a role.
Math has very clear parameters, and requires a direct, analytical thought process. this is also true of chess. And certain studies have shown that playing chess improves mathematical skills in students.
In the game, there are certain principles that all good chess players must know. Time, Point advantage, Space, Development, Pawn Structure, etc.
These all call for three skills: Creativity, circumspension, and MATHEMATIC PROBABILITY. Why? Because all good chess players must look at all the options and possible setbacks of each move. They must think "Ok, I have this many options. If I move here, he has this many options." The goal is to limit the number of options your opponent has.
Also, when it comes to time and developement advantage, you must plan ahead and make sure the number of moves you need to make in order to accomplish your goal is low enough to avoid intervention.
Also, there are simple pawn structures such as a four-pawn chain, doubled up pawns, and deciding whether to start off with one or two spaces.
-A time clock uses math
-Notation requires math
These are all technicalities. But once again, math is not chess's focus. Could you imagine, however, if someone tried to get good at chess without any knowledge of math? Subconciously, our mind takes many mathematical considerations into account without us even thinking about it. For example, there are 64 squares, and I own 16 pieces. This is how much i am winning or losing by, etc.
Chess computers (called engines) use math and probability for positioning. Each piece is assigned a value, the unit of which is the centipawn = 1/100 of a pawn.
Assignments for each piece in centipawns is approximately as follows:
Queen = 900
Rook = 500
Bishop = 300
Knight = 300
Pawn = 100
The King's value is infinite as the game is over once he is captured.
The above values are approximate as competing engines are allowed to assign their own personal value to improve their game.
They are called engines because the software usually does not come with a user interface. User interfaces are written by separate competing companies which often allow the user a choice of which engine to use.
The engine does all the work. It evaluates each legal path for its next move. This is known as a ply. Engines do not evaluate all the way to checkmate as this would take too long, so the interface limits them to an average depth of around 12. Less depth makes the engine easier to beat. More depth requires more computing time. Commercial packages allow the user to set the depth manually.
The engine assigns a value in centipawns to each ply tested, and the ply with the best probability is chosen. This is an oversimplification as the algorithms are sophisticated.
They use probability because most early outcomes are uncertain, and chess games (amongst club level players) are won by tiny cumulative incremental advantages and not by being a more aggressive player.
In general, it is not possible to determine the gender of a child before conception. Gender-specific abortion is illegal in most countries. The only way to be certain of having a male child is to adopt one.
You rolled doubles with a pair of dice what is the probability that the total is less than 9?
You have rolled doubles so the full set of possible outcomes is [1,1], [2,2], [3,3], [4,4],[5,5] and [6,6] - that is six in all.
Of these the first four are favourable outcomes (sum < 9)
So the required probability is 4/6 or 2/3
5 becaese you will get blue
2/9
What is the difference between a discrete and a continuous probability function?
A discrete random variable (RV) can only take a selected number of values whereas a continuous rv can take infinitely many.
Yes. Roughly, very large samples are very likely to have subsets data points having very similar means and distributions. Large numbers of such subsets will tend to be normal distributed (Why?) and will tend to make the total sample be normally distributed.
What is the function of a dandruff?
It has no function. It is simply dead flakes of skin from the scalp.
What are the odds of drawing a fourth ace from a deck of cards if 3 have been dealt?
If three aces have already been dealt with, there is only one ace left and 49 cards left.
P(4th ace)=1/49
The odds are 1 to 49.
It is here assumed that four cards cards are delat one by one and the first three were aces.
What is the role of probability for making inferences in business research?
Examine the role of probability for making inferences in business research.